Grade 12th passMechanics

a particle is influenced by F=yi+zj+xk in a cube of edge length 1m.how can we calculate the workdone along all the paths of that cube

priya
6 Years agoGrade 12th pass

1 Answer

Aastha
6 Years ago
W = f vector dot d vector = k(yi+xj). dr=k (yi+ xj). d (xi+yj) = d(xy) = xy
Now putting limits
W = k(5*7-3*5)=(35-15)k=20kj